OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2). y = 2x + 4 y = negative 1 over 2x − 1 over 2 y = − 1 over 2x − 7 over 2 y = 2x − 8

OpenStudy (anonymous):

@rose21

OpenStudy (jack1):

y = 2x + 4 parallel means must have same gradient so options 2 qnd 3 are out option 1 is the same equation... so its out must b option 4

OpenStudy (jack1):

proof:

OpenStudy (anonymous):

parallel lines have negative answers?

OpenStudy (jack1):

y = 2x − 8 contains point (3, −2)? y = 2x − 8 y = 2*3 − 8 y = 6 − 8 y = -2

OpenStudy (anonymous):

so its D?

OpenStudy (jack1):

parallel lines have negative answers? no, parallel lines have different number for "b" y = mx + b so if equation is identical but with a different B value... it's parallel

OpenStudy (jack1):

is d, but understand y?

OpenStudy (anonymous):

yes thank you so much. how do i give you a medal?

OpenStudy (jack1):

yw, and if u really wanna its the "best response" button ;D

OpenStudy (jack1):

de nada ms fuentes ;D slaters

OpenStudy (anonymous):

ooh okay thanks and i already did <3 may you help me with one more ?

OpenStudy (jack1):

sure, shoot

OpenStudy (anonymous):

Line WX contains (−1, 2) and (4, 12) Line YZ contains points (5, 8) and (−2, −6). Lines WX and YZ are perpendicular because the slopes are the same parallel because the product of the slopes is −1 perpendicular because the product of the slopes is −1 parallel because the slopes are the same

OpenStudy (anonymous):

i used the slope formula for the first two and i got 3.33333

OpenStudy (jack1):

so what is gradient of each...? ("m" value)

OpenStudy (anonymous):

for the second one i got 1.5714

OpenStudy (anonymous):

they are perpendicular

OpenStudy (jack1):

Line WX: (−1, 2) and (4, 12) (x1, y1) and (x2, y2) y2-y1 / x2-x1 =m 12-2 / 4 --1 = m 6/5 =m

OpenStudy (jack1):

Line YZ: (5, 8) and (−2, −6) (x1, y1) and (x2, y2) y2-y1 / x2-x1 =m -6-8 / -2 -5 = m -14/-7 =m

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

for the second one and the first one is 1.2

OpenStudy (anonymous):

oooh now i get it! its D

OpenStudy (anonymous):

they both equal to 2

OpenStudy (jack1):

you're right, Line WX: (−1, 2) and (4, 12) (x1, y1) and (x2, y2) y2-y1 / x2-x1 =m 12-2 / 4 --1 = m 10/5 =m 2 =m sorry... :( a line perpendicular to 2/1 would have a gradient of 1/2 a line parallel to gradient 2 would also have gradient of 2 so they're parallel

OpenStudy (anonymous):

yea thats what i got lol Thank you so much!!!

OpenStudy (jack1):

all good, my bad

OpenStudy (anonymous):

yay! i got a 20/20

OpenStudy (jack1):

props ;D

OpenStudy (anonymous):

you are the best! ill keep you in mind

OpenStudy (jack1):

ooo ;)

OpenStudy (anonymous):

lol byee <3 :*

OpenStudy (jack1):

slaters